#5.The sum of three numbers is 62. The second number is equal to the
first number diminished by 4. The third number is four times the first.
What are the numbers?

Answer by bucky(2189)

Let x represent the first number.
.
The second number equals the first number diminished by 4. So the second number is x - 4.
.
The third number is 4 times the first number. So the third number is 4x.
.
The sum of the three numbers is 62. So we can write the equation:
.
x + x - 4 + 4x = 62
.
Combine the terms in x on the left side (x + x + 4x = 6x) and the equation becomes:
.
6x - 4 = 62
.
Get rid of the -4 on the left side by adding 4 to both sides and you have:
.
6x = 66
.
Solve for x by dividing both sides by 6 and you have:
.
x = 66/6 = 11
.
So the first number is 11. The second number is 11 minus 4 and that is 7. An the third number
is the first number (11) times 4 which is 44.
.
The three numbers are 11, 7, and 44. If you add them up you will find that their total is 62.
.
Hope this helps you to understand the problem.
.